THE DEVELOPMENT AND IMPLEMENTATION OF CHIRPED-PULSE FREQUENCY COMBS AT MIL-LIMETER WAVELENGTHS

\underline{AMANDA L. STEBER}, BRENT J. HARRIS, JUSTIN L. NEILL, KEVIN K. LEHMANN, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904.

Technological advances in such areas as active multiplier chains and high-speed digital electronics are enabling the development of sensitive high-throughput spectroscopic instruments in the millimeter and submillimeter ranges. Recently there has been an effort to develop multiplexed direct absorption spectroscopy techniques that use frequency comb sources derived from phase-locked pulse trains (often created using ultrafast lasers). We have used a high-speed arbitrary waveform generator (AWG with 12 GHz sample rate) to create frequency combs at mm-wave wavelengths using a chirped pulse as the repeating waveform. This waveform has important advantages including 100% duty cycle for the light output and compatibility with the use of frequency multiplier chains that extend the bandwidth of the comb proportional to the frequency multiplication factor. A new spectrometer operating in the 260-290 GHz range using active multiplier chains has been constructed to test the capabilities of chirped-pulse frequency comb spectroscopy for molecular rotational spectroscopy. The spectral properties of the mm-wave combs generated following x24 frequency multiplication, methods for compressed bandwidth detection using a dual-comb approach, and frequency comb analogs of FM spectroscopy will be presented.